(1 point) u(x, t represents the vertical displacement of a string of length L = 20 with wave equation 16. time t = Utt at posshow steps please!
Answer #1

XTC) be solation C letuG pom D 16xTT AgV= XaxoT16AT=0 16T XTCA= o- X)=0 ALO thn xto 4.1bo The &lton a x= G4Q__ The &Pt x=Ge T)=Ge qete-o iRGEOG XC)o olelon Ao B Pthon X4px=o Tre Bole X)t G QDinPX TOG Donppt GGS0+Uro No)-o Dunop = o=ont 20 duTnCan nT n bn=f 20 20 20 mann 20uhth 8 fouie Sine dene in C0,1e 20 uh nb unnTdx 20 Y YUX 40 Casm-o TP n 8odd bn (Uto ud Pom ahae bnas aboleomi 4,)